OpenStudy (anonymous):

THE SUM OF 2 NUMBERS IS 26. AND THEIR POSITIVE DIFFERENCE IS 12. FIND THE LARGER NUMBER

OpenStudy (anonymous):

i want a medal

OpenStudy (anonymous):

OK MEDAL ME AND I MEDAL YOU

OpenStudy (anonymous):

THE SUM OF 2 NUMBERS IS 26. AND THEIR POSITIVE DIFFERENCE IS 12. FIND THE LARGER NUMBER

OpenStudy (anonymous):

HELP

OpenStudy (retireed):

SUM OF 2 NUMBERS IS 26 x+ y =26

OpenStudy (anonymous):

YA

OpenStudy (retireed):

THEIR POSITIVE DIFFERENCE IS 12 translate this english to math

OpenStudy (anonymous):

+12

OpenStudy (retireed):

you forget the DIFFERENCE

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

OMG HELP

OpenStudy (retireed):

Write an equation please for this If I don't see an equation, this will be my last hint THEIR DIFFERENCE IS 12 THEIR refers to X and y

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

CAN I JUST GET THE ANSWER

OpenStudy (retireed):

I really tried hard to help you. Bye.

OpenStudy (anonymous):

WHO WANT'S A FREE MEDAL ?IF YOU DO MEDAL ME AND I WILL MEDAL YOU

OpenStudy (anonymous):

CAN ANYONE HELP ME

OpenStudy (usukidoll):

I'm thinking... hold up.. we are given that the sum of two numbers is 26 x + y = 26 and the difference is 12 difference is like subtraction so maybe if we let y = -12 x-12=26 then solve for x..

OpenStudy (usukidoll):

blah wrong sign

OpenStudy (usukidoll):

x+12=26 now solve for x

OpenStudy (anonymous):

14

OpenStudy (anonymous):

14 IS THE ANSWER ?

OpenStudy (usukidoll):

yeah so the sum of 2 numbers going to 26 is 14+12 = 26 the difference is 12 means that we need to subtract 12 on both sides to get the other number

OpenStudy (anonymous):

BUT THAT IS NOT CHOICE

OpenStudy (anonymous):

14 IS NOT THE ANSWER

OpenStudy (usukidoll):

hmm what are the choices? can you list them out?

OpenStudy (anonymous):

7 15 17 19 21

OpenStudy (anonymous):

QUICKLY REPLY

OpenStudy (usukidoll):

ok let's try this x+y = 26 x-y=12 2x=38 x = 19

OpenStudy (anonymous):

ARE YOU SURE ITS 19 ?

OpenStudy (usukidoll):

seems like there was another equation that we needed to construct -_- 19 + y = 26 y = 7 . I got 19+ 7 = 26 19-7 is the difference of 12 x = 19 and y = 7.. those are the numbers... I'm hoping you can choose those 2

OpenStudy (anonymous):

HOW DID U GET 19

OpenStudy (anonymous):

NVM THX

OpenStudy (anonymous):

FAN AND MEDAL ME

OpenStudy (usukidoll):

I misread the difference is 12 part that's actually another equation x - y =12 so we are supposed to have 2 equations x+y=26 x-y=12 using elimination method we had 2x = 38 and dividing 2 on both sides gives x = 19 and then plug x = 19 back into 19-y=12 to solve for y which was -y=-7 y = 7 since the question was asking for the larger number, it's 19

OpenStudy (anonymous):

YA I KNOW

OpenStudy (anonymous):

I HAVE ANOTHER QUESTION

OpenStudy (usukidoll):

sure

OpenStudy (anonymous):

FIND THE POSITIVE DIFFERENCE BETWEEN X AND Y IF 2x-6=2y+4

OpenStudy (anonymous):

let me post this question